OpenStudy (anonymous):

Does 4x0 = (4x)0? Why or why not?

OpenStudy (anonymous):

Does 4x0 = (4x)0? Why or why not?

OpenStudy (accessdenied):

Like, \(4x^0 = (4x)^0\) ?

OpenStudy (anonymous):

No. its 4x0= (4x)0, why or why not

OpenStudy (accessdenied):

\( 4x \times 0 = (4x)\times 0\) Well, multiplication by 0 in this case always leads to 0. We can order it in any way but it leads to the same inevitable fate -- 0..

OpenStudy (anonymous):

Thanks:)

OpenStudy (accessdenied):

You're welcome! :)